# Fine-Grained Domain Generalization with Feature Structuralization

### v1

@@ III-C Alignment for Commonality, Specificity and Prediction
In essence, our ambition is to endow each feature with a semantic concept that can describe the common and specific parts between objects. In the real world, commonalities and specificities can be identified and utilized to distinguish different objects with the assistance of multi-granularity knowledge. In the same way, this multi-granular structure can effectively assist in identifying and aligning the learned features with those components for DL models.
Commonality Alignment constrains the first section of features to match the common characteristics between objects. The constraint is implemented from two perspectives. Firstly, we argue that the common features within a sample should consistently manifest across diverse granularities because the intrinsic information regarding the objects remains unaffected by alterations in granularities. Secondly, we mandate the congruence of common features belonging to identical categories at the parent granularity but differing at the sub-granularity. For example, zebras and horses at the sub-granularity emerge as distinct sub-categories of parent category Equidae, while sharing certain common attributes, as shown in Fig. 1.
In particular, the output features of $\left.T_{g}\right|_{g=0}^{G-1}$ are initially stacked as a 4-d tensor $F\in\mathbb{R}^{G\times B\times d\times wh}$. All features are tokenized, resulting in $wh$ tokens in dimension $d$, which enables Transformer based variants of FS. The common slice is sequentially extracted from $F$ based on Eq. (1). The first constraint, denoted as $\mathcal{L}_{cs}$, measures how similar the common features of the same samples are across each pair of neighboring granularities. This distance is formulated as
@@
where $B_{k,q,g}$ depicts samples that belong to $q_{th}$ category at granularity $g+1$ and $k_{th}$ sub-category at $g$, respectively. $P_{k,q,g}^{c}$ represents the centroid of the common characteristics of samples belonging to the current sub-granularity category. The second constraint, denoted as $\mathcal{L}_{cd}$, forces the sub-centroids of a given parent category to move closer. The total distance among the stacked sub-centroid tensors $P_{q,g}^{c}\in\mathbb{R}^{K\times d^{c}}$ for all parent categories and all granularities is formulated as
@@
where $Q_{g}$ is the number of parent categories at granularity $g$, and $\mathbf{I}\in\mathbb{R}^{K\times K}$ is a identity matrix.
@@
The total similarity among the stacked centroids $P_{g}^{p}\in\mathbb{R}^{K\times d^{p}}$ for all granularities can be formulated as
$S_{p}=\frac{1}{G-1}\frac{1}{K}\sum_{g=0}^{G-1}SUM({\left\langle P_{g}^{c},P_{g}^{c}\right\rangle\over\|P_{g}^{c}\|_{2}\|P_{g}^{c}\|_{2}}-\mathbf{I}).$ (8)
Prediction Calibration constrains the predictions from each branch in the FSDG model to construct a granularity tree. We impose constraints on multi-granular predictions because solely constraining relationships at the feature level is insufficient. Regarding  [7, 41], the fine-grained ground-truth labels progressively integrate with the predicted distributions at the coarse-grained branches, as shown in Fig. 2. The prediction alignment loss $\mathcal{L}_{lf}$ is designed to connect the prediction spaces of fine and coarse granularities during the training process and to optimize the fine branch. $\mathcal{L}_{lf}$ is formulated as
@@
The FS is achieved by the FS objective function $\mathcal{L}_{FS}$, comprising disentanglement and alignment functions $\mathcal{L}_{dec}$ and $\mathcal{L}_{fa}$, and a prediction alignment loss $\mathcal{L}_{lf}$. $\mathcal{L}_{fa}$ maximizes the similarity of common features while minimizing the similarity of specific features. The FS objective is formulated as
@@ IV Experiments
We train and evaluate the FSDG on three FGDG datasets for image recognition, comparing its generalization performance with several state-of-the-art DG models. Additionally, functional analyses substantiate the effectiveness of FSDG. Codes will be available at https://github.com/YuWLong666/FSDG.
@@ IV-A Datasets
CUB-Paintings consists of two domains, i.e., CUB-200-2011(C) [42] and CUB-200-Paintings(P) [7], with significant domain shifts. Both include four-level granularity, encompassing 14 orders, 38 families, 122 genera, and 200 species. The former C has 11,788 images of real-world bird species, while the latter P is a collection of 3,047 images, including watercolors, oil paintings, pencil drawings, stamps, and cartoons.
@@
TABLE I: Classification accuracy (%) on the Cub-Paintings (RN-50) with best results of DG methods highlighting in bold.
@@ IV-B Implementation and Evaluation
The feature extractor $E$ can be instantiated with ResNet (RN) series [51], Vision Transformer(ViT) series [52], and ASMLP series [53] backbones, respectively. The learning rate is initially set to 0.003 and follows a dynamic strategy, where the coefficient of the learning rate varies from 1 to 0.1 during the training process. All layers except the backbones are trained from scratch, and their learning rate is 10 times that of the backbone layers. The model is optimized using mini-batch SGD with a momentum of 0.9. A single 3090 GPU with a batch size fixed to 32 is used for training and evaluation. Typically, the coefficients are configured as $\lambda_{cs}=0.05$, $\lambda_{cd}=0.5$, and $\lambda_{p}=1$. $r^{c}$ and $r^{p}$ are set to 0.5 and 0.3, respectively. We employ fine-grained classification accuracy as the evaluation criterion. All models are trained three times and evaluated for ten trials on the target domain each time. Due to FSDG retaining only the finest branch during inference, the number of parameters of the dual backbone model is the same as that of the single backbone model, which is 24M.
@@
TABLE II: Classification accuracy (%) on the CompCars (RN-50) with best results of DG methods highlighting in bold.
@@ IV-C Main Results
We compare our method with various DG methods, including ERM [46], ARM [26] , DANN [47], MLDG [27], GroupDRO [48], CORAL [22], SagNet [49], MixStyle [23] , Mixup [24], RIDG [29], SAGM [25], and MIRO [50], based on the DomainBed platform [54] and their official codes. The $\mathcal{L}_{lf}$ in this context is referenced from PAN [7], although it was originally employed in a domain adversarial adaptation scenario rather than for FS in FSDG problems here. We separately present the performance of the baseline FGDG model, FGDG (+$\mathcal{L}_{lf}$) (i.e., replace $\mathcal{L}_{f}$ in Eq. (13) with $\mathcal{L}_{lf}$), and FSDG. We also deploy the FSDG in a single backbone mode (prefixed with S-) for a fairer comparison with other DG methods. The results are shown in Tables  I, II, and III.
On CUB-Paintings, as in Table I, our method performs best across DG competitors. With equivalent training parameter scales, both dual and single backbone FSDGs outperform their respective second-best competitors (i.e., MIRO and SAGM) by 6.64% and 6.86%. In single and dual backbone scenarios, FSDGs exhibit substantial improvements of 2.16% and 1.99% over FGDG (+$\mathcal{L}_{lf}$) models, and 2.58% and 4% over the baselines, respectively. These results validate the superiority of FSDG, which enhances the assistance of coarse to fine granularity and identifies more generalizable fine-grained features. Common features can to some extent represent the invariance of species. They serve as intrinsic elements for classification and are beneficial to generalization. Other DG methods exhibit subpar generalization performance, revealing their vulnerability to fine-grained features.
TABLE III: Classification accuracy (%) on the Birds-31 (RN-50) with best results of DG methods highlighting in bold.
@@
On ComCars, as in Table II, we obtain results with similar trends. FSDGs outperform second-best competitors by 3.07% and 4.59%, surpass baselines by 2.58% and 4.91%, and also outperform FGDG (+$\mathcal{L}_{lf}$) by 1.31% and 1.15%, respectively. The best results even surpass the PAN model. Single backbone models achieve a higher performance than dual backbone models, which can be attributed to the substantial differences between the two domains. It is a challenge to apply multiple levels of clean knowledge learned by dual backbones to a low-quality single-angle domain S when the model is trained on the high-quality multi-angle domain W.
On Birds-31, as in Table III, our method achieves the highest accuracy, surpassing the second-best method, SAGM, by approximately 9.2%. Under the dual backbone configuration, FSDG outperforms the baseline by 2.32%. On average across three datasets, FSDG outperforms the second-best approach by up to 6.2%.
TABLE V: Ablation study on the losses (i.e., $\mathcal{L}_{lf}$, $\mathcal{L}_{cs}$, $\mathcal{L}_{p}$, $\mathcal{L}_{cd}$, and $\mathcal{L}_{dis}$). The classification accuracy (%) is presented.
$\mathcal{L}_{lf}$ $\mathcal{L}_{cs}$ $\mathcal{L}_{p}$ $\mathcal{L}_{cd}$ $\mathcal{L}_{dis}$ C→P P→C Avg
@@
Experiments on various backbones and depths. We conduct experiments on various backbones, including RN-50, RN-101, ViT-Tiny, ViT-Small, ASMLP-Tiny, and ASMLP-Small, as shown in Table IV. Our method exhibits varying degrees of improvement across different frameworks, ranging from 1.9 to 4.6%. Upon replacing the backbone with ViT-S, FSDG achieves a performance of 68.26%, surpassing PAN (DA) by 9.1%, showcasing the superiority of Transformer architecture. FS method achieves the highest performance improvement on ASMLP (4.6%). These results validate the robustness of the FSDG method across various backbone paradigms.
@@ IV-D Analysis
Ablation Study on the losses. We present the outcomes of models optimized under various combinations of losses. As shown in Table V, each loss and its combinations contribute to performance improvement compared to the FGDG baseline (the first row). The efficacy of the FS approach is substantiated by quantitative assessments on FGDG performance.
@@
$S_{cd}$ signifies the distance among sub-centroids’ commonalities within a particular parent category. Fig. 3b shows that $\mathcal{L}_{cd}$ compels an increase in the similarity among different sub-categories. Additionally, we observe a similar effect with $\mathcal{L}_{cs}$. The inclusion of $\mathcal{L}_{cs}$ also increases $S_{cd}$. It can be observed that models lacking alignment tend to perceive common concepts as dissimilar for different sub-classes. The network fails to capture the commonalities among different sub-classes. The reason behind this is that the network without alignment constraints primarily fits discriminative features at a single granularity level, neglecting the shared characteristics among categories. Undoubtedly, this leads to the introduction of spurious discriminative features and a reduction in generalization ability. When both losses are included, $S_{cd}$ reaches its optimal value, highlighting the necessity of employing dual common feature alignment constraints. Overall, to a certain extent, $\mathcal{L}_{cs}$ and $\mathcal{L}_{cd}$ have clustered the sub-categories and learned their common features from the perspective of FS.
@@
TABLE VI: Classification accuracy (%) and Gpu Hours of various distance measure methods. The GPU Hours are tested on conditions: single 3090 GPU, batch size 32, and epoch 1.
@@
TABLE VII: Classification accuracy (%) on various network architectures of the Granularity Transition Layers $T$. BN represents Batch Normalization. Additional configurations for CNN include a stride of 1 and padding of 1.
@@
Experiments on various distance measurements. We conduct experiments with HSIC and Euclidean distance as alternatives of Cosine similarity, as shown in Table. VI. From the perspectives of generalization and computational efficiency, we find that the performance achieved by Cosine exceeds that of other methods by approximately 1%, while the computational time is slightly higher than the Euclidean by 0.01 hours, which is acceptable. Besides, Cosine represents the degree of orthogonality between two vectors which meets our needs. Therefore, we opt to utilize Cosine similarity.
Experiments on various architectures of the Granularity Transition Layer $T$. The Granularity Transition Layer $T$ decouples and refines the globally shared features extracted by the backbone, tailoring them specifically to each granularity classification branch. We configure various structures to construct $T$ with results presented in Table VII. The simplest single-layer CNN achieves the best results. As the number of CNN layers increases, the model’s performance declines, indicating a loss of useful information. Therefore, we choose the single-layer CNN as the Granularity Transition Layer.
@@
Figure 5: Confusion matrix of the concept overlaps Ground Truth and the histogram of the most relevant concepts. (b), (c) and (d) plot the number of top concepts belonging to the decoupled three segments. Avg shows the average value among 8 categories.
@@ V Explainability Analysis
In this paper, we enhance the FGDG performance through the FS method with the integration of data and knowledge. The demonstrated enhancement primarily manifests in that the common part of the features, decoupled based on channel indices, primarily focuses on the commonalities and representing the invariance between categories, while the specific part focuses on learning discriminative characteristics. In this section, we endeavor to validate that FS increases the explicit concept matching intensity between the shared semantic concepts between categories and the indices of model channels.
@@
Figure 6: Confusion matrix of the overlap of relevant concepts between each pair of categories. The first row represents the class relevance of the baseline FGDG model, while the second row belongs to the proposed FSDG method. The first column denotes the overlap of all 26 top relevant channels, the second column denotes those belonging to the common part among the 26 channels, the third column denotes the specific part, and the fourth column denotes the confounding part. The computed confusion matrix is symmetric. To enhance visualization, we set the upper triangle of the confusion matrix to zero. The Ground Truth and the diagonal elements of the confusion matrices are separately plotted in Fig. 5.
@@
We then refine a Concept Relevance Propagation (CRP) technique [18] to acquire the most relevant shared concepts activated by every two categories. Specifically, CRP first records the 40 most relevant samples and their relevance values with respect to each channel of the transition layer $T_{g=0}$. We further refine the CRP by statistically sorting all 256 channels for each class in descending order according to their relevance values. As each channel can represent a concept, this statistical outcome achieves relevance ranking for all 256 concepts with respect to each class. After that, the top 26 channels and their relevance values of each class are extracted to represent the most relevant concepts. The reason why we choose 26 is that, for a single category, only approximately the top 10% of channels are activated, while the remaining 90% have activation levels less than one-tenth of the maximum activation level.
@@
Fig. 5b-5d illustrate the distributions of the top 26 relevant concepts in three parts. We find that FS leads to an increase in the number of concepts activated in the common region while reducing the activation in the other two regions. For example, among the top 26 concepts in category 10, 22 over 26 concepts are activated at the common part by FS, whereas the baseline only has 11. This suggests that, for fine-grained recognition, extracting common concepts is more crucial for generalization as Common concepts to some extent reflect the invariance between species.
@@
Intuitively, we argue that closer categories should share more common concepts, meaning that there should be a higher overlap in the most relevant activated channels. We compute the confusion matrix of the overlap of relevant concepts between each pair of category examples, as illustrated in Fig. 6. Compared among Fig. 5a, 6a and 6e, the overall shape of the concept overlap activated by FS closely resembles the Ground Truth. The FGDG model, without FS, overlooks the relationships between categories, resulting in distant classes having a considerable overlap in activated shared concepts, which is inaccurate. Besides, as shown in Fig. 6b and 6f, FS leads to greater activation of shared concepts in the common part. Besides, the shared concepts between categories are more concentrated in the common partition, which aligns with our initial assumptions and objectives. This indicates that we to some extent pre-defined the functionalities of the indices of the layer channels, enhancing the internal transparency and explainability of the model.
@@
When contrasting the distribution of shared concepts in the specific region, as shown in Fig. 6d and Fig. 6h, we observe that FS elicits a sparser distribution of shared concepts. This indicates that this region focuses more on specificities, aligning with the initial assumptions of the model. The phenomenon of sparse shared concepts also occurs in the confounding part. Since we do not impose constraints on the confounding part, a natural discussion arises as to whether applying constraints only to the common and specific parts would increase more activations in the confounding part. In contrast, by analyzing Fig. 5d, Fig. 6d and Fig. 6h, we find that the number of top activations in the confounding partition is decreased. This implies that FS directs the model to focus more on capturing the relationships between categories, therefore improving generalization.
@@
To quantitatively analyze the matching intensity between shared concepts and channel structure regions, we sum up the number of shared concepts between each category and the other seven categories, as shown in Table VIII. By calculating the proportion of shared concepts belonging to the common region among all shared concepts, we find that FS increases the explicit matching intensity between shared concepts and model channels from 40% to 68%. Therefore, our FS approach is indeed effective.

### v2

@@ III-C Alignment for Commonality, Specificity and Prediction
In essence, our ambition is to endow each feature with a semantic concept that can describe the common and specific parts between objects. In the real world, commonalities and specificities can be identified and utilized to distinguish different objects with the assistance of multi-granularity knowledge. In the same way, this multi-granular structure can effectively assist in identifying and aligning the learned features with those functions for DL models.
Commonality Alignment constrains the first section of features to match the common characteristics between categories. The constraint is implemented from two perspectives. Firstly, we argue that the common features within a sample should consistently manifest across diverse granularities because the intrinsic information regarding the objects remains unaffected by alterations in granularities. Secondly, we mandate the congruence of common features belonging to identical categories at the parent granularity but differing at the sub-granularity. For example, zebras and horses at the sub-granularity emerge as distinct sub-categories of the parent category Equidae while sharing certain common attributes, as shown in Fig. 1.
In particular, the output features of $\left.T_{g}\right|_{g=0}^{G-1}$ are initially stacked as a 4-d tensor $F\in\mathbb{R}^{G\times B\times d\times wh}$. All features are tokenized, resulting in $wh$ tokens in dimension $d$, which enables transformer-based variants of FS. The common slice is sequentially extracted from $F$ based on Eq. (1). The first constraint, denoted as $\mathcal{L}_{cs}$, measures how similar the common features of the same samples are across each pair of neighboring granularities. This distance is formulated as
@@
where $B_{k,q,g}$ depicts samples that belong to $q_{th}$ category at granularity $g+1$ and $k_{th}$ sub-category at $g$. $P_{k,q,g}^{c}$ represents the centroid of the common characteristics of samples belonging to the current sub-granularity category. The second constraint, denoted as $\mathcal{L}_{cd}$, forces the sub-centroids of a given parent category to move closer. The total distance among the stacked sub-centroid tensors $P_{q,g}^{c}\in\mathbb{R}^{K\times d^{c}}$ for all parent categories and all granularities is formulated as
@@
where $Q_{g}$ is the number of parent categories at granularity $g$.
@@
The total distance among the stacked centroids $P_{g}^{p}\in\mathbb{R}^{K\times d^{p}}$ for all granularities can be formulated as
$S_{p}=\frac{1}{G-1}\frac{1}{K}\sum_{g=0}^{G-1}SUM({\left\langle P_{g}^{p},P_{g}^{p}\right\rangle\over\|P_{g}^{p}\|_{2}\|P_{g}^{p}\|_{2}}-\mathbf{I}).$ (8)
Prediction Calibration constrains the predictions from each branch in the FSDG model to construct a granularity tree. Regarding  [7, 41], the fine-grained ground-truth labels progressively integrate with the predicted distributions at the coarse-grained branches, as shown in Fig. 2. The prediction alignment loss $\mathcal{L}_{lf}$ is designed to connect the prediction spaces of fine and coarse granularities during the training process and to optimize the fine branch. $\mathcal{L}_{lf}$ is formulated as
@@
The FS is achieved by the FS objective function $\mathcal{L}_{FS}$, comprising disentanglement and alignment functions $\mathcal{L}_{dec}$ and $\mathcal{L}_{fa}$, and a prediction calibration loss $\mathcal{L}_{lf}$. $\mathcal{L}_{fa}$ maximizes the similarity of common features while minimizing the similarity of specific features. The FS objective is formulated as
@@ IV Experiments
We train and evaluate the FSDG on three FGDG datasets for image recognition, comparing its generalization performance with several state-of-the-art DG methods. Additionally, functional analyses substantiate the effectiveness of FSDG. Codes will be available at https://github.com/YuWLong666/FSDG.
@@ IV-A Datasets
CUB-Paintings consists of two domains, i.e., CUB-200-2011(C) [42] and CUB-200-Paintings(P) [7], with significant domain shifts. Both include four levels of granularity, encompassing 14 orders, 38 families, 122 genera, and 200 species. The former C has 11,788 images of real-world bird species, while P is a collection of 3,047 images, including watercolors, oil paintings, pencil drawings, stamps, and cartoons.
@@
TABLE I: Classification accuracy (%) on the Cub-Paintings (RN-50) with best results of DG methods highlighted in bold.
@@ IV-B Implementation and Evaluation
The feature extractor $E$ can be instantiated with ResNet (RN) series [51], Vision Transformer(ViT) series [52], and ASMLP series [53] backbones, respectively. The learning rate is initially set to 0.003 and follows a dynamic strategy, where the coefficient of the learning rate varies from 1 to 0.1 during the training process. All layers except the backbones are trained from scratch, and their learning rate is 10 times that of the backbone. The model is optimized using mini-batch SGD with a momentum of 0.9. A single 3090 GPU with a batch size fixed to 32 is used for training and evaluation. Typically, the coefficients are configured as $\lambda_{cs}=0.05$, $\lambda_{cd}=0.5$, and $\lambda_{p}=1$. $r^{c}$ and $r^{p}$ are set to 0.5 and 0.3, respectively. We employ fine-grained classification accuracy as the evaluation criterion. All models are trained three times and evaluated for ten trials on the target domain each time. Due to FSDG retaining only the finest branch during inference, the number of parameters of the dual backbone model is the same as that of the single backbone model, which is 24M.
@@
TABLE II: Classification accuracy (%) on the CompCars (RN-50) with best results of DG methods highlighted in bold.
@@ IV-C Main Results
We compare our method with various DG methods, including ERM [46], ARM [26] , DANN [47], MLDG [27], GroupDRO [48], CORAL [22], SagNet [49], MixStyle [23] , Mixup [24], RIDG [29], SAGM [25], and MIRO [50], based on their official codes and the DomainBed platform [54]. The $\mathcal{L}_{lf}$ in this context is referenced from PAN [7], although it was originally employed in a domain adversarial adaptation scenario rather than for FS in FSDG problems here. We separately present the performance of the baseline FGDG model, FGDG (+$\mathcal{L}_{lf}$) (i.e., replace $\mathcal{L}_{f}$ in Eq. (13) with $\mathcal{L}_{lf}$), and FSDG. We also deploy the FSDG in a single backbone mode (prefixed with S-) to compare it with single-backbone DG methods. The results are shown in Tables  I, II, and III.
On CUB-Paintings, as in Table I, our method performs best across DG competitors. With equivalent training parameter scales, both dual and single-backbone FSDGs outperform their respective second-best competitors (i.e., MIRO and SAGM) by 6.64% and 6.86%. In single and dual backbone scenarios, FSDGs exhibit substantial improvements of 2.16% and 1.99% over FGDG (+$\mathcal{L}_{lf}$) models, and 2.58% and 4% over the baselines, respectively. These results validate the superiority of FSDG, which enhances the assistance of coarse to fine granularity and identifies more generalizable fine-grained features. Common features can, to some extent, represent the invariance of species. They serve as intrinsic elements for classification and are beneficial to generalization. Other DG methods exhibit subpar generalization performance, revealing their vulnerability to fine-grained features.
TABLE III: Classification accuracy (%) on the Birds-31 (RN-50) with best results of DG methods highlighted in bold.
@@
On ComCars, as in Table II, we obtain results with similar trends. FSDGs outperform second-best competitors by 3.07% and 4.59%, surpass baselines by 2.58% and 4.91%, and also outperform FGDG (+$\mathcal{L}_{lf}$) by 1.31% and 1.15%, respectively. The best results even surpass the PAN model. Single-backbone models achieve higher performance than dual-backbone models, which can be attributed to the substantial differences between the two domains. It is a challenge to generalize multiple levels of clean knowledge learned by dual backbones to a low-quality single-angle domain S when the model is trained on the high-quality multi-angle domain W.
On Birds-31, as in Table III, our method achieves the highest accuracy, surpassing the second-best method, SAGM, by approximately 9.2%. Under the dual backbone configuration, FSDG outperforms the baseline by 2.32%. On average, across three datasets, FSDG outperforms the second-best approach by up to 6.2%.
TABLE V: Ablation study on the losses (i.e., $\mathcal{L}_{lf}$, $\mathcal{L}_{cs}$, $\mathcal{L}_{p}$, $\mathcal{L}_{cd}$, and $\mathcal{L}_{dec}$). The classification accuracy (%) is presented.
$\mathcal{L}_{lf}$ $\mathcal{L}_{cs}$ $\mathcal{L}_{p}$ $\mathcal{L}_{cd}$ $\mathcal{L}_{dec}$ C→P P→C Avg
@@
Experiments on various backbones and depths. We conduct experiments on various backbones, including RN-50, RN-101, ViT-Tiny, ViT-Small, ASMLP-Tiny, and ASMLP-Small, as shown in Table IV. Our method exhibits varying degrees of improvement across different architectures, ranging from 1.9 to 4.6%. Upon replacing the backbone with ViT-S, FSDG achieves a performance of 68.26%, surpassing PAN (DA) by 9.1%, showcasing the superiority of Transformer architecture. FS method achieves the highest performance improvement on ASMLP (4.6%). These results validate the robustness of the FSDG method across various backbone architectures.
@@ IV-D Analysis
Ablation study on the losses. We present the outcomes of models optimized under various combinations of losses. As shown in Table V, each loss and their combinations contribute to performance improvement compared to the baseline FGDG model (the first row). The efficacy of the FS approach is substantiated by quantitative assessments of FGDG performance.
@@
$S_{cd}$ signifies the distance among sub-centroids’ commonalities within a particular parent category. Fig. 3b shows that $\mathcal{L}_{cd}$ compels an increase in the similarity among different sub-categories. It can be observed that models lacking alignment tend to perceive common concepts as dissimilar. The network fails to capture the commonalities among sub-classes. The reason behind this is that the network without alignment primarily fits discriminative features at a single granularity level, neglecting the shared characteristics among categories. Undoubtedly, this leads to the introduction of spurious discriminative features and a reduction in generalization ability. Additionally, we observe a similar effect with $\mathcal{L}_{cs}$. The inclusion of $\mathcal{L}_{cs}$ also increases $S_{cd}$. When both losses are included, $S_{cd}$ reaches its optimal value, highlighting the necessity of employing dual common feature alignment constraints. Overall, to a certain extent, $\mathcal{L}_{cs}$ and $\mathcal{L}_{cd}$ have clustered the sub-categories and learned their common features from the perspective of FS.
@@
TABLE VI: Classification accuracy (%) and Gpu Hours of various distance metrics. The GPU Hours are tested on conditions: single 3090 GPU, batch size 32, and epoch 1.
@@
TABLE VII: Classification accuracy (%) on various network architectures of the Granularity Transition Layers $T$. BN represents Batch Normalization. Additional configurations for CNN include a stride of 1 and a padding of 1.
@@
Experiments on various distance measurements. We conduct experiments with HSIC and Euclidean distance as alternatives of Cosine similarity, as shown in Table VI. From the perspectives of generalization and computational efficiency, we find that the performance achieved by Cosine exceeds that of other methods by approximately 1%, while the computational time is slightly higher than the Euclidean by 0.01 hours, which is acceptable. Besides, Cosine represents the degree of orthogonality between two vectors, which meets our needs. Therefore, we opt to utilize Cosine similarity.
Experiments on various architectures of the Granularity Transition Layer $T$. $T$ refines the globally shared features extracted by backbones, tailoring them specifically to each granularity branch. We configure various structures to construct $T$ with results presented in Table VII. The simplest single-layer CNN achieves the best results. As the number of CNN layers increases, the model’s performance declines, indicating a loss of useful information. Therefore, we choose the single-layer CNN as the Granularity Transition Layer.
@@
Figure 5: Confusion matrix of the ground truth of concept overlaps and the histograms of the most relevant concepts. (b), (c) and (d) plot the number of top concepts belonging to the three segments. Avg shows the average value among 8 categories.
@@ V Explainability Analysis
In this paper, we enhance the FGDG performance through the FS method with the integration of data and knowledge. The demonstrated enhancement primarily manifests in that the common part of the features, decoupled based on channel indices, primarily focuses on the commonalities and representing the invariance between categories, while the specific part focuses on learning discriminative characteristics. In this section, we endeavor to validate that FS increases the explicit concept matching intensity between the shared concepts among categories and the indices of model channels.
@@
Figure 6: Confusion matrix of the overlap of relevant concepts between each pair of categories. The first row represents the concept overlaps of the baseline FGDG model, while the second row belongs to the proposed FSDG method. The first column denotes the overlap of all 26 top relevant concepts; the second column denotes those belonging to the common part; the third column denotes the specific part; and the fourth column denotes the confounding part. The computed confusion matrix is symmetric. To enhance visualization, we set the upper triangle of the confusion matrix to zero. The ground truth and the diagonal elements of the confusion matrices are separately plotted in Fig. 5.
@@
We then refine a Concept Relevance Propagation (CRP) technique [18] to acquire the most relevant shared concepts activated by every two categories. Specifically, CRP first records the 40 most relevant samples and their relevance values with respect to each channel of the transition layer $T_{g=0}$. We further refine the CRP by statistically sorting all 256 channels for each class according to their relevance values. As each channel can represent a concept, this statistical outcome achieves relevance ranking for all 256 concepts concerning each class. After that, the top 26 channels and their relevance values of each class are extracted to represent the most relevant concepts. The reason why we choose 26 is that, for a single category, only approximately the top 10% of channels are activated, while the remaining 90% have activation levels less than one-tenth of the maximum activation level.
@@
Fig. 5b-5d illustrate the distributions of the top 26 relevant concepts in three parts. We find that FS leads to an increase in the number of concepts activated in the common region while reducing the activation in the other two regions. For example, among the top 26 concepts in category 10, 22 over 26 concepts are activated at the common part by FS, whereas the baseline FGDG model only has 11. This suggests that, for fine-grained recognition, extracting common concepts is more crucial for generalization as common concepts reflect the invariance between categories.
@@
Intuitively, we argue that closer categories should share more common concepts, meaning that there should be a higher overlap in the most relevant activated channels. We compute the confusion matrix of the overlap of relevant concepts between each pair of category examples, as illustrated in Fig. 6. Compared among Fig. 5a, 6a and 6e, the overall shape of the concept overlap activated by FS closely resembles the ground truth. The FGDG model, without FS, overlooks the relationships between categories, resulting in distant classes having a considerable overlap in activated shared concepts, which is inaccurate. Besides, as shown in Fig. 6b and 6f, FS leads to greater activation of shared concepts in the common part. Besides, the shared concepts between categories are more concentrated in the common partition, which aligns with our initial assumptions and objectives. This also indicates that we, to some extent, pre-define the functionalities of the channels, enhancing the model’s internal transparency and explainability.
@@
When contrasting the distribution of shared concepts in the specific region, as shown in Fig. 6d and Fig. 6h, we observe that FS elicits a sparser distribution of shared concepts. This indicates that this region focuses more on specificities, aligning with the initial assumptions of the model. This phenomenon also occurs in the confounding part. Since we do not impose constraints on the confounding part, a natural discussion arises as to whether applying constraints only to the common and specific parts would increase more activations in the confounding part. In contrast, by analyzing Fig. 5d, Fig. 6d and Fig. 6h, we find that the number of top activations in the confounding part is decreased. All of these results demonstrate that FS directs the model to focus more on capturing the relationships between categories rather than on other spurious correlations, therefore improving generalization.
@@
To quantitatively analyze the matching intensity between shared concepts and channel structure regions, we sum up the number of shared concepts between each category and the other seven categories, as shown in Table VIII. By calculating the proportion of shared concepts within the common region relative to all shared concepts, we find that FS increases the explicit matching intensity between shared concepts and model channels from 40% to 68%. Therefore, our FS approach is indeed effective.